OpenStudy (anonymous):

what is the simplified form of 9y^7/15x^5*5x^2/45y^3

OpenStudy (anonymous):

\[\frac{ x^7y^{10} }{ 15 }\]

OpenStudy (anonymous):

(9y^7)/(15x^5) *(5x^2)/(45y^3)

OpenStudy (anonymous):

27y^10 / 5x^7 thats completely wrong

OpenStudy (anonymous):

can you solve it audrey

OpenStudy (anonymous):

hey mertz can you do it

OpenStudy (mertsj):

\[\frac{9y^7}{15x^5}\times\frac{5x^2}{45y^3}=\frac{3y^7}{5x^5}\times\frac{x^2}{9y^3}=\frac{y^4}{15x^3}\]

OpenStudy (anonymous):

thats not one of my answer choices

OpenStudy (mertsj):

What are your answer choices?

OpenStudy (anonymous):

actually thats right thank you so much

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

can you help me with a few more

OpenStudy (mertsj):

Possibly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the simplified form of (y^2-5y+6)/(y^2-2y-3)

OpenStudy (anonymous):

i tried to factor

OpenStudy (mertsj):

\[\frac{y^2-5y+6}{y^2-2y-3}=\frac{(y-3)(y-2)}{(y-3)(y+1)}=\frac{y-2}{y+1}\]

OpenStudy (anonymous):

ok wow im so stupid i guess the brain doesnt work at night

OpenStudy (anonymous):

Thank you and i have a few more

OpenStudy (anonymous):

what is the solution to the rational equation x/2x-1+1/4=2/2x-1

OpenStudy (mertsj):

\[\frac{x}{2x-1}+\frac{1}{4}=\frac{2}{2x-1}\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Multiply each fraction by 4(2x-1). This will cancel each denominator.

OpenStudy (anonymous):

what is the solution to the rational equation x/5+4/5=7

OpenStudy (anonymous):

why do you mutilpy each fraction by 4(2x-1)

OpenStudy (mertsj):

Multiply each term by 5

OpenStudy (mertsj):

To clear the denominators: \[\frac{x}{2x-1}\times4(2x-1)+\frac{1}{4}\times4(2x-1)=\frac{2}{2x-1}\times4(2x-1)\]

OpenStudy (mertsj):

\[4x+2x-1=8\]

OpenStudy (anonymous):

ok thats not one of the answers

OpenStudy (mertsj):

Of course it isn't. You have to finish solving the equation.

OpenStudy (anonymous):

6x-1=8 =x=9/6

OpenStudy (mertsj):

9/6=3/2

OpenStudy (mertsj):

Have to go.

OpenStudy (anonymous):

ok thank you good night